A 120m long train is moving with a velocity of 20m/s and and other of 130m is moving with a velocity of 30m/s in opposite direction.. two trains cross eachother in time.. ?A.. 5secB.. 6secC.. 8sec D.. not

Grade:12th pass

1 Answers

Aabid Hussain
6 years ago
Answer: 5 sec.

Speed of train B with respect to A 50 m/sec.Since train B fully crosses train A , therefore, total distance covered 120+130= 250 m.

Time = Distance/speed

= 250/50 = 5 sec.